Luke borrows $800 000 from a bank to set up a medical practice. He agrees to pay a fixed interest rate of 10.2 per cent per annum (calculated monthly) and to repay by equal monthly instalments over 10 years.

  1. Calculate the monthly repayment.
  2. By how much does Luke’s first repayment reduce the principal?
  3. If the loan is paid off as planned, by how much will the last repayment reduce the principal?
Answer #1

Hello,

The answer are as below

1) Monthly Payment = $10,661

2) Principal payment in First Installment: $3,861

3)   Principal payment in First Installment: $10,571

Please refer to below snapshots for detail solution. Please note, few rows are hidden so that all answers can be see in one frame. All 3 answers are highlighted in yellow.

1 Annual Interest Rate Monthy Interest Rate 10.20% 0.85% Years 10 Efffective No. of Periods PV 120 $800,000 $10,660.86 PMT YeAnnual Interest Rate 0.102 Monthy Interest Rate =C3/12 Years Efffective No. of Periods =C5*12 PV =F3 PMT --PMT(C4,C6,c7) -15